OpenStudy (anonymous):

Help me solve this proportion? x+3 9 ____ = __ 4 x+3

OpenStudy (anonymous):

Do you know cross multiplication? :)

OpenStudy (anonymous):

No. But its multiple choice.

OpenStudy (anonymous):

Okay. So this is how you cross multiply. fraction 1 = fraction 2 product of the numerator of frac 1 and the denominator of frac 2 = product of the deno of frac 1 and numerator of frac 2 from the given: (x+3)(x+3)=(4)(9) (x+3)^2=36 (x+3)^2=6^2 x+3= +or - 6 so, x=6-3 or x=-6-3 got the answer? :)

OpenStudy (anonymous):

x=-6-3 or 6-3

OpenStudy (anonymous):

I got it. Like I get how you do it, but in the multiple choice answers, that's not a choice.

OpenStudy (anonymous):

Not sure if you seen it, but I attached a file that showed the answers and question.

OpenStudy (dumbcow):

haha just add the numbers what is 6-3 ? what is -6-3 ?

OpenStudy (anonymous):

OH! Wow. I'm blonde. SO blonde. hahaha thank you guys!! :)

OpenStudy (dumbcow):

:)